A supervisor has determined that the average salary of the employees in his department $40,000 with a standard deviation of $15,000. Assume that the distribution of the salaries is normally distributed. An employee is selected at random, find the probability that the employee’s salary is:

(i) More than $35,000.
(ii) Between $36,000 to $42,000.
1
Expert's answer
2020-06-23T15:10:37-0400

Let X = the random variable denoting the salary of the selected employee


We have,

X ~ N("\mu" = 40000, "\sigma^2" = 15000)


Then,


Z = "\frac{X-\mu}{\sigma}" ~ N(0, 1), Z is the standard normal variate


(i) The probability that the employee’s salary is more than $35,000


= P(X > 35000)


= P("\frac{X-40000}{15000}>\frac{35000-40000}{15000}")


= P(Z > - 0.33)


= 1 - P(Z "\leq" - 0.33)


= 1 - "\Phi"(- 0.33)


= 1 - 0.3707 [from standard normal table]


= 0.6293


Answer: The probability that the employee’s salary is more than $35,000 is 0.6293.


(ii) The probability that the employee’s salary is between $36,000 to $42,000


P(36000 "\leq" X "\leq" 42000)


= P("\frac{36000-40000}{15000}\leq\frac{X-40000}{15000}\leq\frac{42000-40000}{15000}")


= P(-0.27 "\leq" Z "\leq" 0.13)


= P(Z "\leq" 0.13) - P(Z < -0.27)


= "\Phi"(0.13) - "\Phi"(-0.27)


= 0.5517 - 0.3936 [from standard normal table]


= 0.1581


Answer: The probability that the employee’s salary is between $36,000 to $42,000 is 0.1581.